OpenStudy (kirbykirby):

Consider a sequence of independent trials. At each trial, the possible outcomes are “S” and “F”. On trial n, the probability of getting “S” is equal to Pn = (1/4)^n, n = 1, 2, . . . Let T be the waiting time until the first occurrence of “S”. What is E(T)? (expectation of T)

OpenStudy (anonymous):

i will guess 4

OpenStudy (anonymous):

oh maybe i am misinterpreting the problem let me think before i speak. it looks the probability of getting a success on the nth trial is not \(\frac{1}{4}\) but rather \(\frac{1}{4^n}\)

OpenStudy (kirbykirby):

Yeah.. someone was telling me though expectation infinity, I'm not sure how they got it.

OpenStudy (kirbykirby):

is infinity*

OpenStudy (anonymous):

me neither. i cannot figure out how to add this up

OpenStudy (sirm3d):

when T = 1 (success on the first trial), p = 1/4

OpenStudy (sirm3d):

when T = 2 (success on the 2nd trial), P = (3/4)(1/16) i think the problem iE(T) s the sum of all probabilities.

OpenStudy (anonymous):

probabilities are \[\frac{1}{4}\]\[\frac{3}{4}\times \frac{1}{4^2}\]\[\frac{3}{4}\times \frac{4^2-1}{4^2}\times \frac{1}{4^3}\]\[\frac{3}{4}\times \frac{4^2-1}{4^2}\times \frac{4^3-1}{4^3}\times \frac{1}{4^4}\] etc

OpenStudy (anonymous):

then you have to multiply these by 1,2,3,... and the add my addition is not that good

OpenStudy (kirbykirby):

would this be some kind of geometric distribution?

OpenStudy (anonymous):

if it was i could add it, but it does not look like one to me as there is not common ratio

OpenStudy (anonymous):

the method used for the expected waiting time for bernoulli trials comes from summing the derivative of the geometric series, but this isn't that one either

OpenStudy (kirbykirby):

I'm not too sure what the underlying pmf would be... Can we say though that P(T < infinity) < 1?

OpenStudy (anonymous):

in fact i cant even figure out how to write this one as a formula

OpenStudy (anonymous):

\[P(T=n)=\frac{(4-1)(4^2-1)...(4^{n-1}-1)}{4^{\frac{n(n+1)}{2}}}\] or something like that

OpenStudy (anonymous):

then you have to multiply each term by \(n\) and then you have to sum it up yikes

OpenStudy (kirbykirby):

o.o oh goodness..

OpenStudy (kirbykirby):

this question seemed fairly simple but it is giving me such a headache

OpenStudy (anonymous):

i may be wrong, but i think that is what you have to do on the other hand it is also possible that this series doesn't converge

OpenStudy (anonymous):

no it is not simple lets think for a second maybe the terms don't even go to zero

OpenStudy (anonymous):

well maybe they do sorry i cannot be of more help, maybe re-post and you will get a better answer

OpenStudy (kirbykirby):

Alright I appreciate it the help anyways :)

OpenStudy (kirbykirby):

the*